OpenStudy (anonymous):

What are the roots in the following coordinates? 2x^2-8=0

jimthompson5910 (jim_thompson5910):

2x^2-8=0 2(x^2-4)=0 x^2 - 4 = 0 (x-2)(x+2) = 0 I'll let you finish

OpenStudy (anonymous):

it makes a parabola... positive and negative 2

OpenStudy (anonymous):

can you tell me how to finish it? i dont understand how to do these.

jimthompson5910 (jim_thompson5910):

well if (x-2)(x+2) = 0 then x-2=0 or x+2=0

jimthompson5910 (jim_thompson5910):

solve each of those 2 equations for x

OpenStudy (anonymous):

thank youu.

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

would it be (-2, 0) and (2, 0) or did i do it wrong?

jimthompson5910 (jim_thompson5910):

x = 2 or x = -2 good

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

thanks guys (:

jimthompson5910 (jim_thompson5910):

yw